OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 8x = 34

OpenStudy (anonymous):

(x+4)^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(x+4)^2 is the same as your left hand side + 16 so if you add 16 to both sides you can then write (x+4)^2 = 34+16 = 50

OpenStudy (cunningham2000):

x^2 + 8x = 34 X^2 + 8x + ______ = 34 Divide the 8 by 2 and square it. Add this number to both sides. X^2 + 8x + 16 = 34 + 16